OpenStudy (anonymous):

Can someone please tell me the steps how to graph a quadratic function in intercept form? Here's the problem: f(x)= -4(x+3)(x+7)

OpenStudy (anonymous):

So far i know the x-ints. are (-3,0) (-7,0)

OpenStudy (anonymous):

i think you should change it into a polynomial and go from there but really how i figure these out is with graphing software

OpenStudy (phi):

the average of the two roots is the axis of symmetry

OpenStudy (anonymous):

do i add -3 +(-7 ) over what?? -4?

OpenStudy (phi):

you are finding the middle number between -3 and -7 one way is (-3 + -7)/2 = -10/2 = -5 another way is mark the 2 points and count

OpenStudy (phi):

find y for x=-5. that is where the parabola peaks or bottoms out. Still have to figure out is it a cup or a frown

OpenStudy (anonymous):

Thank you so much this helps a lot , but my math teacher doesnt explain very well all the steps

OpenStudy (phi):

you should get y= -4(-5+3)(-5+7)= -4* -2 * 2 = 16 so it looks like it peaks at (-5,16)

OpenStudy (phi):

as a check, plug in x=0. It should be a large negative number as it zooms down

myininaya (myininaya):